# Jarník-Besicovitch type theorems for semisimple algebraic groups

### v1

@@ 2. Preliminaries
Let $\beta_{0}$ be the weight in the decomposition above such that its weight space $V_{\beta_{0}}$ contains $V_{\lambda_{0}}$. Note that $\beta_{0}$ is defined over $\mathbb{Q}$ and $\beta_{0}=\lambda_{0}|_{\mathbf{T}}$. In particular, $\lambda_{0}(\mathbf{T}(\mathbb{R}))\subset\mathbb{R}$ and $\beta_{0}$ is the highest weight among $\beta$’s relative to $\mathbf{T}$ (the order is determined by the root system $(\Phi,\Phi^{+},\Delta)$) (as discussed in §1).
Since the stable subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\mathbf{P}_{0})$, we have $\alpha(a_{t})\leq 0$ $(t\geq 0)$ for any $\alpha\in\Phi^{+}$. From property (3) of the representation $\rho$ above and the fact that $\textup{Im}\rho\subset\operatorname{\mathbf{SL}}(V)$, one can deduce that $\beta_{0}(a_{t})=\lambda_{0}(a_{t})\leq 0$. If $\beta_{0}(a_{t})=\lambda_{0}(a_{t})=0$, then by the fact that $\textup{Im}\rho\subset\operatorname{\mathbf{SL}}(V)$, for any weight $\lambda$ in $\rho$ relative to $\mathbf{S}$, we have $\lambda(a_{t})=0$ and $\{a_{t}\}_{t\in\mathbb{R}}\subset\ker\rho$, which contradicts the assumption that $\dim\ker\rho=0$. Therefore we have $\beta_{0}(a_{t})=\lambda_{0}(a_{t})<0$ and
@@
Let $\mathbf{P}_{\beta_{0}}$ be the stabilizer of the weight space $V_{\beta_{0}}$ in $\mathbf{G}$. Then $\mathbf{P}_{\beta_{0}}$ contains the minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ since $\beta_{0}$ is the highest weight among the weights $\beta$’s relative to $\mathbf{T}$ in $\rho$. Therefore, $\mathbf{P}_{\beta_{0}}$ is a parabolic $\mathbb{Q}$-subgroup in $\mathbf{G}$. Let $\overline{\mathbf{P}}_{\beta_{0}}$ be the opposite parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{\beta_{0}}$ containing $\overline{\mathbf{P}}_{0}$, and denote by $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ the unipotent radical of $\overline{\mathbf{P}}_{\beta_{0}}$.
Let $N(\mathbf{T})$ and $Z(\mathbf{T})$ be the normalizer and centralizer of $\mathbf{T}$ respectively. Then the Weyl group relative to $\mathbb{Q}$ is defined by
@@
Let $\mathbf{P}_{0}$ be the minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{G}$ as defined above. Then the Bruhat decomposition of $\mathbf{G}$ is the following [6, §21]
$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\mathbf{P}_{0}(\mathbb{Q})$
@@
$\quad\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\overline{\mathbf{P}}_{0}(\mathbb{Q}),\quad\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{Q}).$
If we choose a set $\{w_{i}\}_{i\in I}$ of representatives of ${}_{\mathbb{Q}}{W}$ in $\mathbf{G}(\mathbb{Q})$, then for each $w_{i}$, define
$\mathbf{F}_{w_{i}}=R_{u}(\overline{\mathbf{P}}_{0})\cap w_{i}^{-1}R_{u}(\overline{\mathbf{P}}_{\beta_{0}})w_{i},\quad\mathbf{H}_{w_{i}}=R_{u}(\overline{\mathbf{P}}_{0})\cap w_{i}^{-1}\mathbf{P}_{\beta_{0}}w_{i}$
and we have
$R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{w_{i}}\cdot\mathbf{F}_{w_{i}}.$
Then one can further write the Bruhat decomposition as
$\mathbf{G}(\mathbb{Q})=$ $\bigcup_{i\in I}\mathbf{P}_{0}(\mathbb{Q})\cdot(w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1})(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q}).$
$=$ $\bigcup_{i\in\bar{I}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w_{i}\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})$
where the index sets $I$ and $\bar{I}$ are defined as in §1. We denote by
$\mathcal{C}_{\bar{I}}:=\bigcup_{i\in\bar{I}}w_{i}\mathbf{F}_{w_{i}}(\mathbb{R}).$
Note that the group $w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1}\subset\mathbf{P}_{\beta_{0}}$ is generated by unipotent subgroups whose Lie algebras are direct sums of $\mathbb{Q}$-root spaces in $\mathbf{G}$ relative to $\mathbf{T}$, and $w_{i}\mathbf{H}_{w_{i}}w_{i}^{-1}$ fixes every element in $V_{\beta_{0}}$ by the structure of the representation $\rho$.
@@
An element $g\in\mathbf{G}(\mathbb{R})$ is called rational if $\rho(g)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}$ is Zariski dense in $V_{\beta_{0}}$.
@@
$\rho(\sigma(g)g^{-1})V_{\beta_{0}}=\rho(\sigma(g)g^{-1})\overline{\rho(g)\cdot\Lambda_{g}}=V_{\beta_{0}}$
and $\sigma(g)g^{-1}$ is in the stabilizer $\mathbf{P}_{\beta_{0}}$ of $V_{\beta_{0}}$. Here $\overline{\rho(g)\cdot\Lambda_{g}}$ denotes the Zariski closure of $\rho(g)\cdot\Lambda_{g}$. On the other hand, $\sigma(g)g^{-1}\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})\cap\mathbf{P}_{\beta_{0}}=\{e\}$. Therefore, $\sigma(g)=g$ for any $\sigma\in\operatorname{Gal}(\mathbb{R}/\mathbb{Q})$, and $g\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$. The other direction is clear. ∎
@@
Let $i\in I$ and $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then $w_{i}\cdot g$ is rational if and only if $g\in\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})$.
@@
Let $h_{i}\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $f_{i}\in\mathbf{F}_{w_{i}}(\mathbb{R})$ such that $g=h_{i}\cdot f_{i}$. Suppose that $w_{i}\cdot g$ is rational. Since $w_{i}\in\mathbf{G}(\mathbb{Q})$, by definition, $w_{i}\cdot h_{i}\cdot f_{i}\cdot w_{i}^{-1}$ is also rational in $\mathbf{G}(\mathbb{R})$. We know that $w_{i}\cdot h_{i}\cdot w_{i}^{-1}$ preserves $V_{\beta_{0}}$, so $w_{i}\cdot f_{i}\cdot w_{i}^{-1}$ is rational. By the fact that $w_{i}\mathbf{F}_{w_{i}}w_{i}^{-1}\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and Lemma 2.2, we conclude that $w_{i}\cdot f_{i}\cdot w_{i}^{-1}\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ and $f_{i}\in\mathbf{F}_{w_{i}}(\mathbb{Q})$.
Conversely, suppose that $g=h_{i}\cdot f_{i}$ where $h_{i}\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $f_{i}\in\mathbf{F}_{w_{i}}(\mathbb{Q})$. Then
$w_{i}\cdot g=(w_{i}h_{i}w_{i}^{-1})\cdot(w_{i}f_{i}w_{i}^{-1})\cdot w_{i}.$
Note that $w_{i}h_{i}w_{i}^{-1}\in\mathbf{P}_{\beta_{0}}$ and $w_{i}\in\mathbf{G}(\mathbb{Q})$. By definition, we know that $w_{i}\cdot g$ is rational. This completes the proof of the lemma. ∎
@@
Let $i\in I$ and $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then $w_{i}\cdot g$ is rational if and only if
$g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap(\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma).$
@@
If $w_{i}\cdot g$ is rational, then by Corollary 2.3
$g\in\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{F}_{w_{i}}(\mathbb{Q})\subset\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{G}(\mathbb{Q})=\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{Q})\cdot\mathcal{K}\cdot\Gamma\subset\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma.$
Suppose that $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap(\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma)$. Then there exist $h\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $f\in\mathbf{F}_{w_{i}}(\mathbb{R})$ such that
$g=h\cdot f,\quad f\in\mathbf{H}_{w_{i}}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma.$
@@
$f=h_{i}\cdot p\cdot x\cdot\gamma$
for some $h_{i}\in\mathbf{H}_{w_{i}}(\mathbb{R})$, $p\in\mathbf{P}_{0}(\mathbb{R})$, $x\in\mathcal{K}$ and $\gamma\in\Gamma$. Let $\sigma\in\operatorname{Gal}(\mathbb{R}/\mathbb{Q})$. Then we have
$\sigma(f)=\sigma(h_{i})\cdot\sigma(p)\cdot x\cdot\gamma.$
This implies that $x\cdot\gamma=p^{-1}\cdot h_{i}^{-1}\cdot f=\sigma(p)^{-1}\cdot\sigma(h_{i})^{-1}\cdot\sigma(f)$. Since the product map
$\mathbf{P}_{0}\times\mathbf{H}_{w_{i}}\times\mathbf{F}_{w_{i}}\to\mathbf{G}$
is injective, we have $\sigma(f)=f$ for any $\sigma\in\operatorname{Gal}(\mathbb{R}/\mathbb{Q})$ and hence $f\in\mathbf{F}_{w_{i}}(\mathbb{Q})$. By Corollary 2.3, $w_{i}\cdot g$ is rational. ∎
@@
Let $g$ be a rational element in $\mathbf{G}(\mathbb{R})$. The denominator (or the height) $d(g)$ of $g$ is defined to be the co-volume of the lattice $\rho(g)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}}$ in $V_{\beta_{0}}$.
In the following lemma, we discuss the relation between the shortest vector and the co-volume of a discrete subgroup of the form $\rho(g)\cdot\mathbb{Z}^{n}\cap V_{\beta_{0}}$.
@@
Let $i\in I$ and $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Suppose that $w_{i}\cdot g$ is rational. Then
$d(w_{i}\cdot g)\sim\delta(\rho(w_{i}\cdot g)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}})^{\dim V_{\beta_{0}}}.$
@@
By Corollary 2.3, there exist $h_{i}\in\mathbf{H}_{w_{i}}(\mathbb{R})$ and $f_{i}\in\mathbf{F}_{w_{i}}(\mathbb{Q})$ such that $g=h_{i}\cdot f_{i}$. Note that
$f_{i}\in\mathbf{G}(\mathbb{Q})=w_{i}^{-1}\cdot\mathbf{G}(\mathbb{Q})=w_{i}^{-1}\cdot\mathbf{P}_{0}(\mathbb{Q})\cdot\mathcal{K}\cdot\Gamma$
and there exist $p\in\mathbf{P}_{0}(\mathbb{Q})$, $x\in\mathcal{K}$ and $\gamma\in\Gamma$ such that $f_{i}=w_{i}^{-1}\cdot p\cdot x\cdot\gamma$. We know that $\mathbf{P}_{0}=\mathbf{M}_{0}\cdot R_{u}(\mathbf{P}_{0})$ where $\mathbf{M}_{0}$ is the Levi factor of $\mathbf{P}_{0}$. We write $\mathbf{M}_{a}$ for the maximal $\mathbb{Q}$-anisotropic subgroup in $\mathbf{M}_{0}$. Then $\mathbf{M}_{0}=\mathbf{T}\cdot\mathbf{M}_{a}$, $\mathbf{T}$ commutes with $\mathbf{M}_{a}$ and $\mathbf{M}_{a}(\mathbb{R})/(\mathbf{M}_{a}(\mathbb{Z})\cap\Gamma)$ is compact. So there exist $p_{1}\in\mathbf{T}$, $p_{2}$ in a compact fundamental domain of $\mathbf{M}_{a}(\mathbb{Z})\cap\Gamma$ in $\mathbf{M}_{a}(\mathbb{R})$, $p_{3}\in\mathbf{M}_{a}(\mathbb{Z})\cap\Gamma$ and $u\in R_{u}(\mathbf{P}_{0})$ such that

### v2

@@ 2. Preliminaries
Let $\beta_{0}$ be the weight in the decomposition above such that its weight space $V_{\beta_{0}}$ contains $V_{\lambda_{0}}$. Note that $\beta_{0}$ is defined over $\mathbb{Q}$ and $\beta_{0}=\lambda_{0}|_{\mathbf{T}}$. In particular, $\lambda_{0}(\mathbf{T}(\mathbb{R}))\subset\mathbb{R}$ and $\beta_{0}$ is the highest weight among $\beta$’s relative to $\mathbf{T}$ where the order is determined by the root system $(\Phi,\Phi^{+},\Delta)$.
Since the stable subgroup of $\{a_{t}\}_{t\in\mathbb{R}}$ is contained in $R_{u}(\mathbf{P}_{0})$, we have $\alpha(a_{t})\leq 0$ $(t>0)$ for any $\alpha\in\Phi^{+}$. From property (3) of the representation $\rho$ above and the fact that $\textup{Im}\rho\subset\operatorname{\mathbf{SL}}(V)$, one can deduce that $\beta_{0}(a_{t})=\lambda_{0}(a_{t})\leq 0$ $(t>0)$. If $\beta_{0}(a_{t})=\lambda_{0}(a_{t})=0$ $(t>0)$, then by the fact that $\textup{Im}\rho\subset\operatorname{\mathbf{SL}}(V)$ and $\lambda_{0}$ is the highest weight relative to $\mathbf{S}$, for any other weight $\lambda$ in $\rho$ relative to $\mathbf{S}$, we also have $\lambda(a_{t})=0$ $(t>0)$. This implies that $\{a_{t}\}_{t\in\mathbb{R}}\subset\ker\rho$, which contradicts the assumption that $\dim\ker\rho=0$. Therefore we have $\beta_{0}(a_{t})=\lambda_{0}(a_{t})<0$ $(t>0)$. Note that
@@
Let $\mathbf{P}_{\beta_{0}}$ be the stabilizer of the weight space $V_{\beta_{0}}$ in $\mathbf{G}$. Then $\mathbf{P}_{\beta_{0}}$ contains the minimal parabolic $\mathbb{Q}$-subgroup $\mathbf{P}_{0}$ since $\beta_{0}$ is the highest weight among the weights $\beta$’s relative to $\mathbf{T}$ in $\rho$. Therefore, $\mathbf{P}_{\beta_{0}}$ is a parabolic $\mathbb{Q}$-subgroup in $\mathbf{G}$. It is known that there exists a $\mathbb{Q}$-torus $\mathbf{T}_{\beta_{0}}$ in $\mathbf{T}$ such that $Z(\mathbf{T}_{\beta_{0}})$ (the centralizer of $\mathbf{T}_{\beta_{0}}$ in $\mathbf{G}$) is a Levi subgroup of $\mathbf{P}_{\beta_{0}}$. We denote by $\overline{\mathbf{P}}_{\beta_{0}}$ the opposite parabolic $\mathbb{Q}$-subgroup of $\mathbf{P}_{\beta_{0}}$ containing $\overline{\mathbf{P}}_{0}$ with the same Levi subgroup $Z(\mathbf{T}_{\beta_{0}})$, and denote by $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ the unipotent radical of $\overline{\mathbf{P}}_{\beta_{0}}$.
Let $N(\mathbf{T})$ and $Z(\mathbf{T})$ be the normalizer and centralizer of $\mathbf{T}$ in $\mathbf{G}$ respectively. Then the Weyl group relative to $\mathbb{Q}$ is defined by
@@
Let $\mathbf{P}_{0}$ be the minimal parabolic $\mathbb{Q}$-subgroup of $\mathbf{G}$ as defined above, and $_{\mathbb{Q}}\mathcal{W}$} a set of representatives of $_{\mathbb{Q}}{W}$ in $N(\mathbf{T})(\mathbb{Q})$. Then the Bruhat decomposition of $\mathbf{G}$ is the following [8, Theorem 21.15]
$\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}{W}}\cdot\mathbf{P}_{0}(\mathbb{Q}).$
Note that the Weyl group $_{\mathbb{Q}}{W}$ acts transitively on the set of minimal parabolic $\mathbb{Q}$-subgroups containing $Z(\mathbf{T})$ [8, Corollary 21.3]. So there exists an element in $_{\mathbb{Q}}{W}$ which sends $\mathbf{P}_{0}$ to $\overline{\mathbf{P}}_{0}$, and we let $\bar{w}$ be the representative of this element in $_{\mathbb{Q}}\mathcal{W}$}. Then we have
$\mathbf{G}(\mathbb{Q})=\mathbf{G}(\mathbb{Q})\cdot\bar{w}^{-1}=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}\mathcal{W}}\cdot(\mathbf{P}_{0}(\mathbb{Q})\cdot\bar{w}^{-1})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}\mathcal{W}}\cdot\overline{\mathbf{P}}_{0}(\mathbb{Q})$
@@
$\quad\mathbf{G}(\mathbb{Q})=\mathbf{P}_{0}(\mathbb{Q})\cdot{{}_{\mathbb{Q}}\mathcal{W}}\cdot R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{Q}).$
For each $w\in{{}_{\mathbb{Q}}\mathcal{W}}$, define
$\mathbf{F}_{w}=R_{u}(\overline{\mathbf{P}}_{0})\cap w^{-1}R_{u}(\overline{\mathbf{P}}_{\beta_{0}})w,\quad\mathbf{H}_{w}=R_{u}(\overline{\mathbf{P}}_{0})\cap w^{-1}\mathbf{P}_{\beta_{0}}w$
and by [8, Proposition 21.9], we have
$R_{u}(\overline{\mathbf{P}}_{0})=\mathbf{H}_{w}\cdot\mathbf{F}_{w}.$
We may assume that $e\in{{}_{\mathbb{Q}}\mathcal{W}}$. It follows that
$\mathbf{G}(\mathbb{Q})=$ $\bigcup_{w\in{{}_{\mathbb{Q}}\mathcal{W}}}\mathbf{P}_{0}(\mathbb{Q})\cdot(w\mathbf{H}_{w}w^{-1})(\mathbb{Q})\cdot w\cdot\mathbf{F}_{w}(\mathbb{Q})$
$=$ $\bigcup_{w\in{{}_{\mathbb{Q}}\mathcal{W}}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w\cdot\mathbf{F}_{w}(\mathbb{Q}).$
Here, as discussed in §1.2, the subsets $\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w\cdot\mathbf{F}_{w}(\mathbb{Q})$ $(w\in{{}_{\mathbb{Q}}\mathcal{W}})$ may overlap, and we choose any subset $_{\mathbb{Q}}\overline{\mathcal{W}}$} in $_{\mathbb{Q}}\mathcal{W}$} (as small as possible) such that
$\mathbf{G}(\mathbb{Q})=\bigcup_{w\in{{}_{\mathbb{Q}}\overline{\mathcal{W}}}}\mathbf{P}_{\beta_{0}}(\mathbb{Q})\cdot w\cdot\mathbf{F}_{w}(\mathbb{Q}).$
Note that, by [8, Proposition 21.9], the group $w\mathbf{H}_{w}w^{-1}\subset\mathbf{P}_{\beta_{0}}$ is generated by unipotent subgroups whose Lie algebras are sums of root spaces of positive $\mathbb{Q}$-roots in $\mathfrak{g}$ relative to $\mathbf{T}$, and $\beta_{0}$ is the highest weight in $\rho$. Hence $w\mathbf{H}_{w}w^{-1}$ fixes every element in $V_{\beta_{0}}$ by the structure of the representation $\rho$.
Now we prove Theorem 1.3.
Proof of Theorem 1.3.
Since $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is bounded, by definition
$\tau(\psi)=\liminf_{t\to\infty}\left(-\frac{\ln(\psi(t))}{t}\right)\geq 0.$
Note that we have already proved $\beta_{0}(a_{-1})>0$. Now suppose that $\tau=\tau(\psi)>\beta_{0}(a_{-1})$. Then for any sufficiently small $\epsilon>0$, there exists $C_{\epsilon}>0$ such that
$\psi(t)\leq C_{\epsilon}\cdot e^{-(\tau-\epsilon)t}\quad(\forall t>0).$
Let $g\in\mathbf{G}(\mathbb{R})$. Since $\beta_{0}(a_{-1})>0$ is the fastest contracting rate of the $\{a_{t}\}_{t\in\mathbb{R}}$-action on $V$, we have
$\|\rho(a_{t})v\|\geq e^{-\beta_{0}(a_{-1})t}\|v\|\quad(\forall v\in\rho(g)\cdot\mathbb{Z}^{d}\setminus\{0\})$
and
$\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})\geq e^{-\beta_{0}(a_{-1})t}\cdot\delta(\rho(g)\cdot\mathbb{Z}^{d}).$
Since $\tau>\beta_{0}(a_{-1})$, by choosing $\epsilon$ sufficiently small, one may find a constant $C>0$ such that
$\delta(\rho(a_{t}\cdot g)\cdot\mathbb{Z}^{d})\geq C\cdot\psi(t)\quad(t>0).$
which shows that $g\in S_{\rho}(\psi)$ and $S_{\rho}(\psi)=\mathbf{G}(\mathbb{R})$. Note that this argument also works for $\tau=\infty$. This completes the proof of Theorem 1.3. ∎
In the rest of the paper, we assume that $0\leq\tau(\psi)\leq\beta_{0}(a_{-1})$. In the following, we define the notion of rational elements and discuss some related properties.
@@
An element $g\in\mathbf{G}(\mathbb{R})$ is called rational if $\rho(g)\mathbb{Z}^{d}\cap V_{\beta_{0}}$ is Zariski dense in $V_{\beta_{0}}$, or equivalently, if $\rho(g)\mathbb{Z}^{d}\cap V_{\beta_{0}}$ is a lattice in the vector space of real points in $V_{\beta_{0}}$.
@@
$\rho(\sigma(g)g^{-1})V_{\beta_{0}}=\rho(\sigma(g)g^{-1})\overline{\rho(g)\cdot\Lambda_{g}}=\overline{\rho(\sigma(g))\cdot\Lambda_{g}}=V_{\beta_{0}}$
and $\sigma(g)g^{-1}$ is in the stabilizer $\mathbf{P}_{\beta_{0}}$ of $V_{\beta_{0}}$. Here $\overline{\rho(g)\cdot\Lambda_{g}}$ and $\overline{\rho(\sigma(g))\cdot\Lambda_{g}}$ denote the Zariski closures of $\rho(g)\cdot\Lambda_{g}$ and $\rho(\sigma(g))\cdot\Lambda_{g}$ respectively. On the other hand, $\sigma(g)g^{-1}\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and $R_{u}(\overline{\mathbf{P}}_{\beta_{0}})\cap\mathbf{P}_{\beta_{0}}=\{e\}$. Therefore, $\sigma(g)=g$ for any $\sigma\in\operatorname{Gal}(\mathbb{R}/\mathbb{Q})$, and $g\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$. The other direction is clear. ∎
@@
Let $w\in{{}_{\mathbb{Q}}\mathcal{W}}$ and $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then $w\cdot g$ is rational if and only if $g\in\mathbf{H}_{w}(\mathbb{R})\cdot\mathbf{F}_{w}(\mathbb{Q})$.
@@
Let $h\in\mathbf{H}_{w}(\mathbb{R})$ and $f\in\mathbf{F}_{w}(\mathbb{R})$ such that $g=h\cdot f$. Suppose that $w\cdot g$ is rational. Since $w\in\mathbf{G}(\mathbb{Q})$, by definition, $w\cdot g\cdot w^{-1}=w\cdot h\cdot f\cdot w^{-1}$ is also rational in $\mathbf{G}(\mathbb{R})$. Since $w\cdot h\cdot w^{-1}$ preserves $V_{\beta_{0}}$, we get that $w\cdot f\cdot w^{-1}$ is rational. By the fact that $w\mathbf{F}_{w}w^{-1}\subset R_{u}(\overline{\mathbf{P}}_{\beta_{0}})$ and Lemma 2.2, we conclude that $w\cdot f\cdot w^{-1}\in R_{u}(\overline{\mathbf{P}}_{\beta_{0}})(\mathbb{Q})$ and $f\in\mathbf{F}_{w}(\mathbb{Q})$.
Conversely, suppose that $g=h\cdot f$ where $h\in\mathbf{H}_{w}(\mathbb{R})$ and $f\in\mathbf{F}_{w}(\mathbb{Q})$. Then
$w\cdot g=(whw^{-1})\cdot(wfw^{-1})\cdot w.$
Note that $whw^{-1}\in\mathbf{P}_{\beta_{0}}$ and $w\in\mathbf{G}(\mathbb{Q})$. By definition, we know that $w\cdot g$ is rational. This completes the proof of the corollary. ∎
@@
Let $w\in{{}_{\mathbb{Q}}\mathcal{W}}$ and $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Then $w\cdot g$ is rational if and only if
$g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap(\mathbf{H}_{w}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma).$
@@
If $w\cdot g$ is rational, then by Corollary 2.3, we have
$g\in\mathbf{H}_{w}(\mathbb{R})\cdot\mathbf{F}_{w}(\mathbb{Q})\subset\mathbf{H}_{w}(\mathbb{R})\cdot\mathbf{G}(\mathbb{Q})=\mathbf{H}_{w}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{Q})\cdot\mathcal{K}\cdot\Gamma\subset\mathbf{H}_{w}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma.$
Now suppose that $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})\cap(\mathbf{H}_{w}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma)$. Then there exist $h\in\mathbf{H}_{w}(\mathbb{R})$ and $f\in\mathbf{F}_{w}(\mathbb{R})$ such that
$g=h\cdot f,\quad f\in\mathbf{H}_{w}(\mathbb{R})\cdot\mathbf{P}_{0}(\mathbb{R})\cdot\mathcal{K}\cdot\Gamma.$
@@
$f=h\cdot p\cdot x\cdot\gamma$
for some $h\in\mathbf{H}_{w}(\mathbb{R})$, $p\in\mathbf{P}_{0}(\mathbb{R})$, $x\in\mathcal{K}$ and $\gamma\in\Gamma$. Let $\sigma\in\operatorname{Gal}(\mathbb{R}/\mathbb{Q})$. Then we have
$\sigma(f)=\sigma(h)\cdot\sigma(p)\cdot x\cdot\gamma.$
This implies that $x\cdot\gamma=p^{-1}\cdot h^{-1}\cdot f=\sigma(p)^{-1}\cdot\sigma(h_{i})^{-1}\cdot\sigma(f)$. Since the product map
$\mathbf{P}_{0}\times\mathbf{H}_{w}\times\mathbf{F}_{w}\to\mathbf{G}$
is injective, we have $\sigma(f)=f$ for any $\sigma\in\operatorname{Gal}(\mathbb{R}/\mathbb{Q})$ and hence $f\in\mathbf{F}_{w}(\mathbb{Q})$. By Corollary 2.3, $w\cdot g$ is rational. ∎
@@
Let $g$ be a rational element in $\mathbf{G}(\mathbb{R})$. The height $0pt(g)$ of $g$ is defined to be the co-volume of the lattice $\rho(g)\mathbb{Z}^{d}\cap V_{\beta_{0}}$ in the vector space of real points in $V_{\beta_{0}}$.
In the following lemma, we discuss the relation between the length of the shortest non-zero vector and the co-volume of the discrete subgroup $\rho(g)\cdot\mathbb{Z}^{n}\cap V_{\beta_{0}}$ for $g\in\mathbf{G}(\mathbb{R})$. For convenience, we write $A\lesssim B$ $(A\gtrsim B)$ if there exists a constant $c>0$ such that
$A\leq c\cdot B\quad(A\geq c\cdot B).$
If $A\lesssim B$ and $A\gtrsim B$, then we write $A\asymp B$. We will specify the implicit constants in the contexts if necessary.
@@
Let $w\in{{}_{\mathbb{Q}}\mathcal{W}}$ and $g\in R_{u}(\overline{\mathbf{P}}_{0})(\mathbb{R})$. Suppose that $w\cdot g$ is rational. Then
$0pt(w\cdot g)\asymp\delta(\rho(w\cdot g)\cdot\mathbb{Z}^{d}\cap V_{\beta_{0}})^{\dim V_{\beta_{0}}}.$
@@
By Corollary 2.3, there exist $h\in\mathbf{H}_{w}(\mathbb{R})$ and $f\in\mathbf{F}_{w}(\mathbb{Q})$ such that $g=h\cdot f$. Note that
$f\in\mathbf{G}(\mathbb{Q})=w^{-1}\cdot\mathbf{G}(\mathbb{Q})=w^{-1}\cdot\mathbf{P}_{0}(\mathbb{Q})\cdot\mathcal{K}\cdot\Gamma$
and there exist $p\in\mathbf{P}_{0}(\mathbb{Q})$, $x\in\mathcal{K}$ and $\gamma\in\Gamma$ such that $f=w^{-1}\cdot p\cdot x\cdot\gamma$. We know that $\mathbf{P}_{0}=\mathbf{M}_{0}\cdot R_{u}(\mathbf{P}_{0})$ where $\mathbf{M}_{0}$ is the Levi factor of $\mathbf{P}_{0}$. We write $\mathbf{M}_{a}$ for the maximal $\mathbb{Q}$-anisotropic subgroup in $\mathbf{M}_{0}$. Then $\mathbf{M}_{0}=\mathbf{T}\cdot\mathbf{M}_{a}$, $\mathbf{T}$ commutes with $\mathbf{M}_{a}$ and $\mathbf{M}_{a}(\mathbb{R})/(\mathbf{M}_{a}(\mathbb{Z})\cap\Gamma)$ is compact. So there exist $p_{1}\in\mathbf{T}$, $p_{2}$ in a compact fundamental domain of $\mathbf{M}_{a}(\mathbb{Z})\cap\Gamma$ in $\mathbf{M}_{a}(\mathbb{R})$, $p_{3}\in\mathbf{M}_{a}(\mathbb{Z})\cap\Gamma$ and $u\in R_{u}(\mathbf{P}_{0})$ such that